OpenStudy (anonymous):

Find the equation in standard form which passes through (3, –6) and (–2, –1). You must show all work. I am so stuck... :(

OpenStudy (anonymous):

First lets find the slope. You know how?

OpenStudy (anonymous):

y2 - y1/x2 - x1??

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

so whats the slope

OpenStudy (anonymous):

5/-5

OpenStudy (anonymous):

great..thats -1

OpenStudy (anonymous):

so do you know the eqn of a line... slope = (y-y1)/(x-x1) ??

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I plug it into that now?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

is it just y - -6 for the top?

OpenStudy (anonymous):

precisely

OpenStudy (anonymous):

Ah, i'm confused

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

(y - - 6) / (x-3) = -1 wts wrong?

OpenStudy (anonymous):

didn't we just do that?

OpenStudy (anonymous):

thats all simplify that (y+6)/(x-3) = -1 so y+6 = 3-x

OpenStudy (anonymous):

it should be y - y1 = m(x - x1)

OpenStudy (anonymous):

same thing

OpenStudy (anonymous):

isnt it (y- - 6) = -1(x-3)?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

i figured it out

OpenStudy (anonymous):

so the answer is?

OpenStudy (anonymous):

y = -x + 6

OpenStudy (anonymous):

no

OpenStudy (anonymous):

-1 + 6 which is 5 so y = 5

OpenStudy (anonymous):

y + 6 = 3 - x x+y + 3 =0

OpenStudy (anonymous):

isnt it?

OpenStudy (anonymous):

I don't know, i just typed in what i just said... I'm just stuck on this other one now

OpenStudy (anonymous):

ok..lets see that